MPD filter development..
Source: http://2012portal.blogspot.com/2020/05/blog-post_13.html
MPD filter development..
Source: http://2012portal.blogspot.com/2020/05/blog-post_13.html
VTX stabilized, basic filtration protocol established..
Source: http://2012portal.blogspot.com/2020/05/blog-post_1.html
Pandora in progress, ongoing gamma grid ratio disruption, L2LSIP temporary abort, Project 502 on hold. GR+. Drastic VTX / Systems security breach at 504, deflected. Minimum M / HP requirements met, VTX / SD / DL requirements not met…
Source: http://2012portal.blogspot.com/2020/05/state-of-mission-report-april-2020.html
URGENT EXIT protocol fully enforced..
Source: http://2012portal.blogspot.com/2020/04/blog-post_25.html
Top domain parameter secure, evaporation sequence..
Source: http://2012portal.blogspot.com/2020/04/blog-post_10.html
Pandora in progress, ongoing gamma grid ratio disruption, PBHL TOP anomaly detected. L2LSIP unstable, Project 502 on hold. Minimum M / DL / HP requirements met, VTX and SD requirements not met. Significant M insertion…
Source: http://2012portal.blogspot.com/2020/04/state-of-mission-report-march-2020.html
Pandora in progress, gamma grid ratio disruption. L2LSIP stable, Project 502 operational. Minimum M / DL / HP requirements met, VTX and SD requirements not met…
Source: http://2012portal.blogspot.com/2020/03/state-of-mission-report-february-2020.html
L2LSIP operational, stability test..
Source: http://2012portal.blogspot.com/2020/02/blog-post_5.html
Minimum L2LSIP security requirements met, modulation in progress, P502v1…
Source: http://2012portal.blogspot.com/2020/02/blog-post_25.html
Initialize Project 502, modulation target 200223…
Source: http://2012portal.blogspot.com/2020/02/initialize-project-502-modulation.html
L4 Pandora sequence in progress. GR+, exit optimization in progress. Minimum M / DL / HP requirements met, VTX and SD requirements not met…
Source: http://2012portal.blogspot.com/2020/02/state-of-mission-report-january-2020.html
L4 Pandora sequence in progress, project 501 merged into MASTERPIECE. VTXC systems security breach deflected. Minimum M / DL / HP requirements met, VTX and SD requirements not met…
Source: http://2012portal.blogspot.com/2019/12/state-of-mission-report-november-2019.html
MOSS Pandora sequence in progress, project 501 to merge into MASTERPIECE at ST point. VTXC systems security breach, in deflection, brief RM2m L0 MASTERPIECE sequence prevented L0 collapse. Minimum M / DL / HP requirements met, VTX and SD requirements not met…
Source: http://2012portal.blogspot.com/2019/11/state-of-mission-report-october-2019.html
All protocols clear and operational, 501/MASTERPIECE integration set at ST point. INT YXR flow recovery almost complete, except for recent VTXC disruption which modifies YXR protocols (VTX.1 downgrade). Project 310 aborted and project 061 downgraded due to unfavorable grid ratio…
Source: http://2012portal.blogspot.com/2019/09/project-501-update.html
On the cosmic scale, the shift into the next cosmic cycle continues…
EXMOSS Pandora sequence in progress. Project 310 total abort due to security breach, systems security breach recovery in progress. Minimum M / HP / VTX requirements met, SD requirements not met, DL requirements not met…
Source: http://2012portal.blogspot.com/2019/09/state-of-mission-report-august-2019.html
Security breach at 310, project permanently aborted..
Source: http://2012portal.blogspot.com/2019/08/blog-post_28.html
Security breach deflected, sublunar EELA in position..
Source: http://2012portal.blogspot.com/2019/07/blog-post_23.html
EXMOSS Pandora sequence in progress, cosmic rootkit detected, removal in progress, N22 detected, removal in progress. Project 501 final integration into MASTERPIECE in progress, internal YXR flow / multiphasic recovery in progress, PB recovery failure…
Source: http://2012portal.blogspot.com/2019/07/state-of-mission-report-june-2019.html
Systems alert downgrade to yellow..
Source: http://2012portal.blogspot.com/2019/06/blog-post_23.html